# SynCo: Synthetic Hard Negatives in Contrastive Learning for Better Unsupervised Visual Representations

### v3

@@ Abstract
Contrastive learning has become a dominant approach in self-supervised visual representation learning. Hard negatives - samples closely resembling the anchor - are key to enhancing learned representations’ discriminative power. However, efficiently leveraging hard negatives remains challenging. We introduce SynCo (Synthetic negatives in Contrastive learning), a novel approach that improves model performance by generating synthetic hard negatives on the representation space. Building on the MoCo framework, SynCo introduces six strategies for creating diverse synthetic hard negatives on-the-fly with minimal computational overhead. SynCo achieves faster training and better representation learning, reaching 67.9% top-1 accuracy on ImageNet ILSVRC-201 linear evaluation after 200 pretraining epochs, surpassing MoCo’s 67.5% using the same ResNet-50 encoder. It also transfers more effectively to detection tasks: on PASCAL VOC, it outperforms both the supervised baseline and MoCo with 82.6% AP; on COCO, it sets new benchmarks with 41.0% AP for bounding box detection and 35.7% AP for instance segmentation. Our synthetic hard negative generation approach significantly enhances visual representations learned through self-supervised contrastive learning. Code is available at https://github.com/giakoumoglou/synco.
@@ 1 Introduction
Figure 1: SynCo extends MoCo by introducing synthetic hard negatives generated on-the-fly from a memory queue. The process begins with two augmented views of an image, $\mathbf{x}_{q}$ and $\mathbf{x}_{k}$, processed by an encoder and a momentum encoder, respectively, producing feature vectors $\mathbf{q}$ and $\mathbf{k}$. The memory queue holds negative samples $\mathbf{n}_{1},\mathbf{n}_{2},\ldots$, which are concatenated with synthetic hard negatives $\mathbf{h}_{1},\mathbf{h}_{2},\ldots$ generated using the SynCo strategies. These combined negatives are used to compute the affinity matrix, which, together with the positive pair (query $\mathbf{q}$ and key $\mathbf{k}$), contributes to the InfoNCE loss calculation.
@@
    We empirically show improved downstream performance on ImageNet by incorporating synthetic hard negatives, demonstrating enhancements in both linear evaluation and semi-supervised learning tasks.
@@
    We show SynCo learns stronger representations by measuring their transfer learning capabilities COCO and PASCAL VOC detection, where it outperforms both the supervised baseline and MoCo.
@@ 2.2 Hard Negatives
Hard negatives are critical in contrastive learning as they improve the quality of visual representations by helping to define the representation space more effectively. These challenging yet relevant samples are harder to distinguish from the anchor point, enabling the model to better differentiate between similar features. The use of hard negatives involves selecting samples that are similar to positive samples but different enough to aid in learning distinctive features. Dynamic sampling of hard negatives during training prevents the model from easily minimizing the loss, enhancing its learning capabilities [20, 6]. Various approaches have been proposed to leverage hard negatives effectively. For instance, MoCo [20] utilizes a dynamic queue and momentum-based encoder updates to maintain fresh and challenging negatives throughout training. Other methods, such as SimCLR [6] and InfoMin [39], suggest adjusting the difficulty of negative samples by varying data augmentation techniques. This progressive increase in task difficulty benefits the training process. Building on these ideas, MoCHI [23] has explored integrating hard negative mixing into existing frameworks to further improve performance. When incorporating these approaches into representation space generation, it’s essential to consider how negative samples impact the model’s learning space. The goal is to implement a strategy that not only challenges the model but also enhances its ability to extract and differentiate features effectively. By employing these methods, models become more adept at handling detailed and complex tasks, ensuring each negative sample significantly contributes to optimizing learning outcomes and boosting overall model effectiveness.
@@ 2.3 Synthetic Features
Synthetic feature generation is a widely used method to enhance deep learning models, especially in cases with limited labeled data. By adding synthetic features to the representation space, models improve in generalization and performance. Methods have generated features for unseen classes using generative models [17, 48, 36], while others have integrated these into self-supervised and contrastive learning frameworks [27, 54]. Models that combine generation with representation spaces have also shown success in zero-shot learning [16]. In contrast, our approach directly generates synthetic hard negatives in contrastive learning, improving representation without additional generative models.
@@ 3.2 Understanding Hard Negatives
Hardness of negatives. The ”hardness” of negative samples, defined by their similarity to positive samples in the representation space, determines how challenging they are for the model to differentiate, directly impacting the effectiveness of the contrastive learning process. Figure 2 illustrates the evolution of negative sample hardness during MoCo-v2 training on ImageNet-100. The plot depicts the top 1024 matching probabilities $p_{z_{i}}$ across different training epochs. Initially, the distribution of these probabilities is relatively uniform. However, as training progresses, a clear trend emerges: fewer negatives contribute significantly to the loss function. This observation suggests that the model rapidly learns to distinguish most negatives, leaving only a small subset that remains challenging. Such a phenomenon underscores the importance of maintaining a diverse pool of hard negatives throughout the training process to sustain effective learning [23].
@@
Difficulty of the proxy task. The difficulty of the proxy task in contrastive learning, typically defined by the self-supervised objective, significantly influences the quality of learned representations. Figure 3 compares the proxy task performance of MoCo and MoCo-v2 on ImageNet-100, measured by the percentage of queries where the key ranks above all negatives. Notably, MoCo-v2, which employs more aggressive augmentations, exhibits lower proxy task performance compared to MoCo, indicating a more challenging learning objective. Paradoxically, this increased difficulty correlates with improved performance on downstream tasks such as linear classification [23]. This counterintuitive relationship between proxy task difficulty and downstream performance suggests that more challenging self-supervised objectives can lead to the learning of more robust and transferable representations, motivating the development of strategies to dynamically modulate task difficulty during training.
@@ 4 Synthetic Hard Negatives in Contrastive Learning
Based on these observations, we present an approach for generating synthetic hard negatives in the representation space using six distinct strategies. We refer to the proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
4.1 Hard Negatives Generation
@@
Interpolated hard negatives (type 1). For each query $\mathbf{q}$, we propose to generate $N_{1}$ synthetic hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{1}=\{\mathbf{h}_{1}^{1},\mathbf{h}_{2}^{1},\ldots,\mathbf{h}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{1}\in H^{1}$ would be given by:
$\mathbf{h}_{k}^{1}=\alpha_{k}\cdot\mathbf{q}+(1-\alpha_{k})\cdot\mathbf{n}_{i},\quad\alpha_{k}\in(0,\alpha_{\text{max}})$ (4)
@@
Extrapolated hard negatives (type 2). For each query $\mathbf{q}$, we propose to generate $N_{2}$ hard negative features by extrapolating beyond the query embedding in the direction of the hardest negative features. Similar to the interpolated method, we use a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{2}=\{\mathbf{h}_{1}^{2},\mathbf{h}_{2}^{2},\ldots,\mathbf{h}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{2}\in H^{2}$ would be given by:
$\mathbf{h}_{k}^{2}=\mathbf{n}_{i}+\beta_{k}\cdot(\mathbf{n}_{i}-\mathbf{q}),\quad\beta_{k}\in(1,\beta_{\text{max}})$ (5)
@@
Mixup hard negatives (type 3). For each query $\mathbf{q}$, we propose to generate $N_{3}$ hard negative features by combining pairs of the $N$ hardest existing negative features in $\mathcal{\hat{Q}}^{N}$. Let $H^{3}=\{\mathbf{h}_{1}^{3},\mathbf{h}_{2}^{3},\ldots,\mathbf{h}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{3}\in H^{3}$ would be given by:
$\mathbf{h}_{k}^{3}=\gamma_{k}\cdot\mathbf{n}_{i}+(1-\gamma_{k})\cdot\mathbf{n}_{j},\quad\gamma_{k}\in(0,1)$ (6)
@@
Noise-injected hard negatives (type 4). For each query $\mathbf{q}$, we propose to generate $N_{4}$ hard negative features by adding Gaussian noise to the hardest negative features. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{4}=\{\mathbf{h}_{1}^{4},\mathbf{h}_{2}^{4},\ldots,\mathbf{h}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{4}\in H^{4}$ would be given by:
$\mathbf{h}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ (7)
@@
Perturbed hard negatives (type 5). For each query $\mathbf{q}$, we propose to generate $N_{5}$ hard negative features by perturbing the embeddings of the hardest negative features. Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{5}=\{\mathbf{h}_{1}^{5},\mathbf{h}_{2}^{5},\ldots,\mathbf{h}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{5}\in H^{5}$ would be given by:
$\mathbf{h}_{k}^{5}=\mathbf{n}_{i}+\delta\cdot\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i})$ (8)
@@
Adversarial hard negatives (type 6). For each query $\mathbf{q}$, we propose to generate $N_{6}$ hard negative features by applying adversarial perturbations to the hardest negative features to maximize their similarity to the query embeddings. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{6}=\{\mathbf{h}_{1}^{6},\mathbf{h}_{2}^{6},\ldots,\mathbf{h}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{h}_{k}^{6}\in H^{6}$ would be given by:
$\mathbf{h}_{k}^{6}=\mathbf{n}_{i}+\eta\cdot\text{sign}(\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i}))$ (9)
@@ 4.2 Discussion on Synthetic Hard Negatives
Intuition of synthetic hard negatives. Each SynCo strategy imrpoves model generalization through challenging contrasts. Type 1 interpolates between query and hard negatives, increasing sample diversity throughout training. Type 2 extrapolates beyond the query, pushing representation space boundaries and improving robustness to difficult contrasts. Type 3 combines pairs of hard negatives, encouraging more generalized and robust feature learning. Type 4 injects Gaussian noise, promoting invariance to minor feature fluctuations and enhancing generalization. Type 5 modifies embeddings based on similarity gradients, refining discriminatory power by directing the model towards harder negatives. Type 6 applies adversarial perturbations, creating the most challenging contrasts to distinguish deceptively similar samples. A simplified example illustrating the synthetic hard negative generation is provided in Figure 4. The latter figure visualizes the different types of synthetic hard negatives generated from a query and three original negative features in a 2D representation space. Dotted lines connect the query to extrapolated negatives and between pairs of mixup negatives, illustrating the relationships used to generate these synthetic negative features.
Figure 4: A toy example illustrating the six types of synthetic hard negatives generated by SynCo from a query and three original negative features in a 2D representation space.
Is the proxy task more difficult? Figure 3 depicts the proxy task performance for different configurations of SynCo. We observe that incorporating synthetic negatives leads to faster learning and improved performance. Each type of synthetic negative accelerates learning compared to the MoCo-v2 baseline, with the full SynCo configuration showing the most significant improvement (see Figure 6(a)) and the lowest final proxy task performance. This indicates that SynCo presents the most challenging proxy task. This is evidenced by $\max\ell(\mathbf{h}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, where $\mathbf{h}_{k}^{i}\in H^{i}$ are synthetic negatives and $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$ are original negatives. Through SynCo, we modulate proxy task difficulty via synthetic negatives, pushing the model to learn more robust features.
Evaluating the usage of the representation space. To assess learned representations, we employ alignment and uniformity metrics proposed by [44]. These metrics provide insights into representation space utilization, with alignment quantifying the grouping of similar samples and uniformity measuring representation spread across the hypersphere. Figure 5 presents results for various models using features from the ImageNet-100 validation set. Our findings demonstrate that SynCo significantly enhances the uniformity of representations compared to MoCo-v2 and MoCHI, indicating its superior ability to utilize the representation space in the proxy task. Furthermore, the incorporation of synthetic negatives (types 1 to 6) leads to improved alignment. These results suggest that SynCo’s approach to synthetic negative generation and contrastive learning yields more effective and well-distributed feature representations.

### v4

@@ Abstract
Contrastive learning has become a dominant approach in self-supervised visual representation learning. Hard negatives - samples closely resembling the anchor - are key to enhancing learned representations’ discriminative power. However, efficiently leveraging hard negatives remains challenging. We introduce SynCo (Synthetic negatives in Contrastive learning), a novel approach that improves model performance by generating synthetic hard negatives on the representation space. Building on the MoCo framework, SynCo introduces six strategies for creating diverse synthetic hard negatives on-the-fly with minimal computational overhead. SynCo achieves faster training and better representation learning, reaching 67.9% top-1 accuracy on ImageNet ILSVRC-2012 linear evaluation after 200 pretraining epochs, surpassing MoCo’s 67.5% using the same ResNet-50 encoder. It also transfers more effectively to detection tasks: on PASCAL VOC, it outperforms both the supervised baseline and MoCo with 82.5% AP; on COCO, it sets new benchmarks with 40.9% AP for bounding box detection and 35.5% AP for instance segmentation. Our synthetic hard negative generation approach significantly enhances visual representations learned through self-supervised contrastive learning. Code is available at https://github.com/giakoumoglou/synco.
@@ 1 Introduction
Figure 1: SynCo extends MoCo [21, 9] by introducing synthetic hard negatives generated on-the-fly from a memory queue. The process begins with two augmented views of an image, $\mathbf{x}_{q}$ and $\mathbf{x}_{k}$, processed by an encoder and a momentum encoder, respectively, producing feature vectors $\mathbf{q}$ and $\mathbf{k}$. The memory queue holds negative samples $\mathbf{n}_{1},\mathbf{n}_{2},\ldots$, which are concatenated with synthetic hard negatives $\mathbf{s}_{1},\mathbf{s}_{2},\ldots$ generated using the SynCo strategies. These combined negatives are used to compute the affinity matrix, which, together with the positive pair (query $\mathbf{q}$ and key $\mathbf{k}$), contributes to the InfoNCE loss calculation.
@@
    We empirically show improved downstream performance on ImageNet ILSVRC-2012 by incorporating synthetic hard negatives, demonstrating improvements in both linear evaluation and semi-supervised learning tasks.
@@
    We show that SynCo learns stronger representations by measuring their transfer learning capabilities COCO and PASCAL VOC detection, where it outperforms both the supervised baseline and MoCo.
@@ 2.2 Hard Negatives
Hard negatives are critical in contrastive learning as they improve the quality of visual representations by helping to define the representation space more effectively. These challenging yet relevant samples are harder to distinguish from the anchor point, enabling the model to better differentiate between similar features. The use of hard negatives involves selecting samples that are similar to positive samples but different enough to aid in learning distinctive features. Dynamic sampling of hard negatives during training prevents the model from easily minimizing the loss, enhancing its learning capabilities [21, 7]. Various approaches have been proposed to leverage hard negatives effectively. For instance, MoCo [21] utilizes a dynamic queue and momentum-based encoder updates to maintain fresh and challenging negatives throughout training. Other methods, such as SimCLR [7] and InfoMin [40], suggest adjusting the difficulty of negative samples by varying data augmentation techniques. This progressive increase in task difficulty benefits the training process. Building on these ideas, MoCHI [24] has explored integrating hard negative mixing into existing frameworks to further improve performance. By employing these methods, models become more adept at handling detailed and complex tasks, ensuring each negative sample significantly contributes to optimizing learning outcomes and boosting overall model effectiveness.
@@ 2.3 Synthetic Features
Synthetic feature generation is a widely used method to enhance deep learning models, especially in cases with limited labeled data. By adding synthetic features to the representation space, models improve in generalization and performance. Some methods have generated features for unseen classes using generative models [18, 49, 37], while others have integrated these into self-supervised and contrastive learning frameworks [28, 55]. Models that combine generation with representation spaces have also shown success in zero-shot learning [17]. In contrast, our approach directly generates synthetic hard negatives in contrastive learning, improving representation without additional generative models.
@@ 3.2 Understanding Hard Negatives
Hardness of negatives.
The ”hardness” of negative samples, defined by their similarity to positive samples in the representation space, determines how challenging they are for the model to differentiate, directly impacting the effectiveness of the contrastive learning process. Figure 2 illustrates the evolution of negative sample hardness during MoCo-v2 training on ImageNet-100. The plot depicts the top 1024 matching probabilities $p_{z_{i}}$ across different training epochs. Initially, the distribution of these probabilities is relatively uniform. However, as training progresses, a clear trend emerges: fewer negatives contribute significantly to the loss function. This observation suggests that the model rapidly learns to distinguish most negatives, leaving only a small subset that remains challenging. Such a phenomenon underscores the importance of maintaining a diverse pool of hard negatives throughout the training process to sustain effective learning [24].
@@
Difficulty of the proxy task.
The difficulty of the proxy task in contrastive learning, typically defined by the self-supervised objective, significantly influences the quality of learned representations. Figure 3 compares the proxy task performance of MoCo and MoCo-v2 on ImageNet-100, measured by the percentage of queries where the key ranks above all negatives. Notably, MoCo-v2, which employs more aggressive augmentations, exhibits lower proxy task performance compared to MoCo, indicating a more challenging learning objective. Paradoxically, this increased difficulty correlates with improved performance on downstream tasks such as linear classification [24]. This counterintuitive relationship between proxy task difficulty and downstream performance suggests that more challenging self-supervised objectives can lead to the learning of more robust and transferable representations, motivating the development of strategies to dynamically modulate task difficulty during training.
@@ 4 Synthetic Hard Negatives in Contrastive Learning
In this section, we present an approach for generating synthetic hard negatives in the representation space using six distinct strategies. We refer to our proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
4.1 Generating Synthetic Hard Negatives
@@
Interpolated synthetic negatives (type 1).
For each query $\mathbf{q}$, we propose to generate $N_{1}$ synthetic hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $S^{1}=\{\mathbf{s}_{1}^{1},\mathbf{s}_{2}^{1},\ldots,\mathbf{s}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{s}_{k}^{1}\in S^{1}$ would be given by:
$\mathbf{s}_{k}^{1}=\alpha_{k}\cdot\mathbf{q}+(1-\alpha_{k})\cdot\mathbf{n}_{i},\quad\alpha_{k}\in(0,\alpha_{\text{max}})$ (4)
@@
Extrapolated synthetic negatives (type 2).
For each query $\mathbf{q}$, we propose to generate $N_{2}$ hard negative features by extrapolating beyond the query embedding in the direction of the hardest negative features. Similar to the interpolated method, we use a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $S^{2}=\{\mathbf{s}_{1}^{2},\mathbf{s}_{2}^{2},\ldots,\mathbf{s}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{s}_{k}^{2}\in S^{2}$ would be given by:
$\mathbf{s}_{k}^{2}=\mathbf{n}_{i}+\beta_{k}\cdot(\mathbf{n}_{i}-\mathbf{q}),\quad\beta_{k}\in(1,\beta_{\text{max}})$ (5)
@@
Mixup synthetic negatives (type 3).
For each query $\mathbf{q}$, we propose to generate $N_{3}$ hard negative features by combining pairs of the $N$ hardest existing negative features in $\mathcal{\hat{Q}}^{N}$. Let $S^{3}=\{\mathbf{s}_{1}^{3},\mathbf{s}_{2}^{3},\ldots,\mathbf{s}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{s}_{k}^{3}\in S^{3}$ would be given by:
$\mathbf{s}_{k}^{3}=\gamma_{k}\cdot\mathbf{n}_{i}+(1-\gamma_{k})\cdot\mathbf{n}_{j},\quad\gamma_{k}\in(0,1)$ (6)
@@
Noise-injected synthetic negatives (type 4).
For each query $\mathbf{q}$, we propose to generate $N_{4}$ hard negative features by adding Gaussian noise to the hardest negative features. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $S^{4}=\{\mathbf{s}_{1}^{4},\mathbf{s}_{2}^{4},\ldots,\mathbf{s}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{s}_{k}^{4}\in S^{4}$ would be given by:
$\mathbf{s}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ (7)
@@
Perturbed synthetic negatives (type 5).
For each query $\mathbf{q}$, we propose to generate $N_{5}$ hard negative features by perturbing the embeddings of the hardest negative features. Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $S^{5}=\{\mathbf{s}_{1}^{5},\mathbf{s}_{2}^{5},\ldots,\mathbf{s}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{s}_{k}^{5}\in S^{5}$ would be given by:
$\mathbf{s}_{k}^{5}=\mathbf{n}_{i}+\delta\cdot\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i})$ (8)
@@
Adversarial synthetic negatives (type 6).
For each query $\mathbf{q}$, we propose to generate $N_{6}$ hard negative features by applying adversarial perturbations to the hardest negative features to maximize their similarity to the query embeddings. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $S^{6}=\{\mathbf{s}_{1}^{6},\mathbf{s}_{2}^{6},\ldots,\mathbf{s}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated. Then a synthetic negative feature $\mathbf{s}_{k}^{6}\in S^{6}$ would be given by:
$\mathbf{s}_{k}^{6}=\mathbf{n}_{i}+\eta\cdot\text{sign}(\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i}))$ (9)
@@
4.2 Integrating Synthetic Hard Negatives into the Contrastive Loss
The synthetic hard negatives generated are integrated into the contrastive learning process by modifying the InfoNCE loss. Let $\mathcal{S}=\bigcup_{i=1}^{6}S^{i}$ represent the concatenation of all synthetic hard negatives, where $S^{i}$ is the set of synthetic negatives generated by the $i$-th strategy. This combined set of synthetic negatives augments the original negatives $\mathcal{Q}$, providing a more diverse and challenging set of contrasts for the query. The modified InfoNCE loss is given by:
$\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q},\mathcal{S})=-\log\frac{\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)}{Z+\sum_{\mathbf{s}\in\mathcal{S}}\exp(\mathbf{q}^{T}\cdot\mathbf{s}/\tau)}$ (10)
where $Z=\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)+\sum_{\mathbf{n}\in\mathcal{Q}}\exp(\mathbf{q}^{T}\cdot\mathbf{n}/\tau)$, $\tau$ is the temperature parameter, $\mathcal{Q}$ is the set of original negatives, and $\mathcal{S}$ is the set of synthetic hard negatives. By incorporating both real and synthetic negatives, the model is exposed to a wider variety of challenging examples, which encourages learning more robust and generalizable representations. The overall computational overhead of SynCo is roughly equivalent to increasing the queue/memory by $\sum_{i=1}^{6}N_{i}\ll K$, along with the additional cost of generating the synthetic negatives.
@@ 4.3 Discussion on Synthetic Hard Negatives
The following explores the effects of incorporating synthetic hard negatives on the difficulty of the proxy task and how they influence the usage of the representation space.
Is the proxy task more difficult?
Figure 3 depicts the proxy task performance for different configurations of SynCo. We observe that incorporating synthetic negatives leads to faster learning and improved performance. Each type of synthetic negative accelerates learning compared to the MoCo-v2 baseline, with the full SynCo configuration showing the most significant improvement (see Figure 5(a)) and the lowest final proxy task performance. This indicates that SynCo presents the most challenging proxy task. This is evidenced by $\max\ell(\mathbf{s}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, where $\mathbf{s}_{k}^{i}\in S^{i}$ are synthetic negatives and $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$ are original negatives. Through SynCo, we modulate proxy task difficulty via synthetic negatives, pushing the model to learn more robust features.
Evaluating the usage of the representation space.
To assess learned representations, we employ alignment and uniformity metrics proposed by [45]. These metrics provide insights into representation space utilization, with alignment quantifying the grouping of similar samples and uniformity measuring representation spread across the hypersphere. Figure 4 presents results for various models using features from the ImageNet-100 validation set. Our findings demonstrate that SynCo significantly enhances the uniformity of representations compared to MoCo-v2 and MoCHI, indicating its superior ability to utilize the representation space in the proxy task. Furthermore, the incorporation of synthetic negatives (types 1 to 6) leads to improved alignment. These results suggest that SynCo’s approach to synthetic negative generation and contrastive learning yields more effective and well-distributed feature representations.